In an experiment to determine the specific charge (e/m) of an electron, electrons where accelerated by an electric field strength of 10N/C and the magnetic field strength of 0.1H is used to deflect them. Find the speed of the accelerated electrons. What is the radius of the path described by the electron?
1
Expert's answer
2021-02-04T02:34:27-0500

a) Using the Newton’s Second Law of Motion we can write:


"evB=\dfrac{mv^2}{r},""r=\dfrac{mv}{eB}."

Let's assume that both fields are adjusted so that remains undeflected and arrives at the point at the center of the screen. Then, we can write:


"F_e=F_m,""Bev=eE,""v=\dfrac{E}{B}=\dfrac{E}{\mu_0H}=\dfrac{10\ \dfrac{N}{C}}{4\pi\cdot10^{-7}\ \dfrac{H}{m}\cdot0.1\ H}=7.96\cdot10^7\ \dfrac{m}{s}."

Then, we can find the radius of the path described by the electron:


"r=\dfrac{9.1\cdot10^{-31}\ kg\cdot7.96\cdot10^7\ \dfrac{m}{s}}{1.6\cdot10^{-19}\ C\cdot4\pi\cdot10^{-7}\ \dfrac{H}{m}\cdot0.1\ H}=3602\ m."